A LOT IS IN THE SHAPE OF THE RIGHT TRIANGLE. THE SHORTER LEG MEASURES 180 METERS. THE HYPOTENUSE IS 60 METERS LONGER THAN THE LENGTH OF THE LONGER LEG. HOW LONG IS THE LONGER LEG?
A)240m
B)300m
C)360m
D)180m

Answer by nerdybill(7384)

A LOT IS IN THE SHAPE OF THE RIGHT TRIANGLE. THE SHORTER LEG MEASURES 180 METERS. THE HYPOTENUSE IS 60 METERS LONGER THAN THE LENGTH OF THE LONGER LEG. HOW LONG IS THE LONGER LEG?
A)240m
B)300m
C)360m
D)180m
.
Apply Pythagorean theorem:
Let x = length of longer leg
then
180 = shorter leg
60+x = hypotenuse
.
x^2 + 180^2 = (60+x)^2
x^2 + 180^2 = (60+x)(60+x)
x^2 + 32400 = 3600 + 120x + x^2
32400 = 3600 + 120x
32400 = 3600 + 120x
28800 = 120x
240 meters = x